Project

General

Profile

02012017

Nightly Observing Plan for MJD 57786 (starting evening of Feb 1, 2017)

Half night

Observers

  • Alex Drlica-Wagner (Obs 1)
  • Ting Li (Obs 2, RM)

CTIO Staff

  • Hernán Tirado - Obs. Support
  • Alberto Alvarez - Operator
  • Staff calendar

Almanac

mjd 57786, date 20170201, Site CTIO, UT-offset 3 hr
Current local time 14:25 CLST     UT 17:25     SidTim 21:31
Sunset     20:49 CLST  23:49 UT   SidTim   59 deg
Eve. -10   21:26 CLST  00:26 UT   SidTim   68 deg  Zeropoint; standards
Eve. -14   21:47 CLST  00:47 UT   SidTim   74 deg  Begin obstac
Eve. -18   22:09 CLST  01:09 UT   SidTim   79 deg  Last chance standards
Moonset    23:51 CLST  02:51 UT   SidTim  104 deg
Midpoint   01:57 CLST  04:57 UT   SidTim  136 deg
Morn -18   05:44 CLST  08:44 UT   SidTim  193 deg  Begin astronomical twilight
Morn Std   06:06 CLST  09:06 UT   SidTim  198 deg  End obstac; begin standards
Morn -10   06:27 CLST  09:27 UT   SidTim  203 deg
Sunrise    07:05 CLST  10:05 UT   SidTim  213 deg
Moon position (deg): RA 17.1, Dec 4.1 Phase: 0.326

Weather Conditions

Observing Plan and Goals for the Night

* Evening Standards: Yzirg.
UT Date = '2017/02/02 00:26'     Local time = 21:26
E4-A                  ra 09:23:44  dec -45:21:02  secz 1.9
SDSSJ0200-0000        ra 02:00:00  dec +00:00:00  secz 1.5
SDSSJ0320-0000        ra 03:20:00  dec +00:00:00  secz 1.2

* Defocus script
* Obstac: Mostly wide field
* Cosmos field around 1:15 CT (4:15 UT)
  • r-band deep script if seeing in r < 1.0",
  • If not, then g-band deep if g<1.1",
  • If not, then i-band deep if i<1.0",
  • If not, then short field zY if i < 1.05" .

* Half Night standards (using calspec) use only 1 (highest airmass): grizY

UT Date = '2017/02/02 04:52'     Local time = 01:52
SDSSJ1227-0000        ra 12:27:00  dec +00:00:00  secz 1.9

When were the last SN fields observed?

(See https://cdcvs.fnal.gov/redmine/projects/desops/wiki/When_were_the_SN_fields_last_observed for database query.)

hex_name     seqtype     last_complete
SN-X3    long_z    2017-01-26 00:55:51.528351+00:00
SN-E1    short    2017-01-27 00:53:18.591071+00:00
SN-X2    short    2017-01-27 01:14:13.977544+00:00
SN-X3    long_i    2017-01-27 01:33:32.157988+00:00
SN-E2    short    2017-01-28 00:53:06.852121+00:00
SN-X1    short    2017-01-28 01:14:15.665466+00:00
SN-X3    long_r    2017-01-28 01:33:29.425048+00:00
SN-X3    long_g    2017-01-29 01:04:14.492806+00:00
SN-S2    short    2017-01-29 01:17:20.589302+00:00
SN-S1    short    2017-01-29 01:36:41.706064+00:00
SN-C3    long_r    2017-01-29 01:57:12.786249+00:00
SN-C3    long_z    2017-01-29 02:20:30.702561+00:00
SN-C2    short    2017-01-31 00:48:51.012551+00:00
SN-C3    long_g    2017-01-31 01:12:08.842422+00:00
SN-C3    long_i    2017-01-31 01:33:23.812696+00:00
SN-C1    short    2017-01-31 02:22:33.431244+00:00

Notes from 4pm meeting

Run Manager's Meeting Feb. 01, 2017
-----------------------------------
HTD, Mandeep, Eric N, Bobby B, Alex, Ting, Hernan, Michael, Huan, David F.
Brian N., Robert

Camera & Telescope
------------------
No problems, flats in a few minutes.

Last Night
----------
http://des-ops.fnal.gov:8080/nightsum/nightsum-2017-01-31/nightsum.html
Start of observing Observer2 crashed. At start of Observer2, have to restart SISPI.
Came back with a bad Pan. restarted the pan and reconfigured. Then things worked.
Lost 30 minutes. Skipped standards.

Then WF with good seeing. For last 40 minutes did the 2nd half of gri Cosmos field.

exposure 615119 maybe not delivered - maybe doesn't exist - needs to be declared bad by Eric N.
615120 looked weird. In one backplane, backplane 3, the background seems smaller.
for now we'll call this OK and have a closer look.
An earthquake at 03:07 UT. Pointing wasn't affected. But we should look for double features.
Maybe the earthquake knocked out the DIMM.

There was a tiling 6 exposure yesterday: 615076. That's weird; we thought 6 was done.

DESDM & DQ
-----------
Got all the data. All WF good.

We've done the gri short Cosmos Fields. Still have the zY to do.

Forecast
--------
Weather looks good. Overhead clear.
SN deadman is off.

Huan will provide deep cosmos field, 6 sons of about 30 minutes.
Tonight do an r-band script if past < 1.0",
If not, then g-band if better than 1.1",
Then if not, then i-band if better 1.0",
Then if not, then short field zY if better than 1.05" in i-band.

AOB
----
Rasicam indicator for Blanco position wasn't working correctly.
Shift is to see what is happening tonight, looking at the Rasicam computer.
WoD: These aren't the stars you're looking for.

Wildlife report: Zorro’s, no puma\

Translating PSF's between band:
if i = 1.0",
then g = 1.10", r=1.04", z=0.96", Y=0.95 independent of airmass.

airmass corrections: airmass ratio to 3/5 power.